# Text to Band Gap: Pre-trained Language Models as Encoders for Semiconductor Band Gap Prediction

### v2

@@ 3.2 Model Performance
We evaluated the performance of three transformer-based language models, RoBERTa, T5, and LLaMA-3, on the task of predicting semiconductor band gaps from text-based material descriptions. Each model was finetuned with a custom regression head and assessed using two types of input formats: a structured string representation and a natural language description generated via GPT-3.5 Turbo. These models were compared against shallow ML baselines, which also use the same input data format), including support vector regression (SVR), random forest, and XGBoost. The accuracy of the model was quantified using three metrics: The mean absolute error (MAE), the root mean square error (RMSE), and the coefficient of determination ($R^{2}$), as shown in Table 2, with the parity plots visualized in Figure 2.
Table 2: Comparison of model performance across different ML approaches. Each transformer-based model was evaluated using both structured string and description input formats. Best performance per metric is shown in bold.
@@
Figure 2: Parity plots for band gap predictions across models: a SVR, b XGBoost, c Random Forest, d RoBERTa, e T5, f LLaMA-3
Across all models, inputs in structured string format consistently outperformed descriptive natural language inputs. Among the transformer-based models, LLaMA-3 with structured input achieved the best overall performance, with a mean absolute error of 0.248 eV and R² of 0.891. While RoBERTa and T5 also outperformed the shallow ML models, their performance was lower when using GPT-generated descriptions compared to structured string inputs. However, the ability of all three language models to still learn meaningful correlations from unstructured natural language input underscores a key strength of this approach. Rather than viewing this variability as a limitation, we interpret it as evidence of the models’ capacity to generalize from flexible, human-like descriptions of materials. Notably, even in its smallest size, LLaMA-3 demonstrated strong capacity for property prediction when paired with fine-tuning, indicating that decoder-only architectures can serve as lightweight yet powerful alternatives for scientific regression tasks.
@@ 3.3 Layer Freezing Analysis
To assess the extent to which each model’s pretrained representations contribute to band gap prediction, we conducted layer freezing experiments for RoBERTa, T5, and LLAMA-3 using structured string inputs. In each experiment, we progressively froze layers and finetuned only a portion of the model. The goal was to determine how much domain-specific tuning is necessary to achieve high predictive accuracy. Table 3 report the MAE, RMSE, and $R^{2}$ values for each model under various freezing configurations. In all cases, a custom regression head was trained on top of the partially or fully frozen transformer backbone.
Table 3: Comparison of layer freezing strategies across RoBERTa, T5, and LLaMA-3 using structured string inputs. The fully finetuned (non-frozen) results are included for reference. All values are reported as mean $\pm$ standard deviation.
Model Freezing Strategy MAE (eV) RMSE (eV) $R^{2}$
RoBERTa Fully finetuned (no freezing) $0.325\pm 0.006$ $0.447\pm 0.005$ $0.855\pm 0.004$
Freeze first layer $0.328\pm 0.008$ $0.448\pm 0.006$ $0.848\pm 0.005$
Freeze all but final 3 layers $0.388\pm 0.009$ $0.510\pm 0.007$ $0.817\pm 0.006$
Freeze all but final layer $0.509\pm 0.012$ $0.648\pm 0.009$ $0.721\pm 0.009$
Freeze all layers $0.663\pm 0.016$ $0.826\pm 0.011$ $0.563\pm 0.013$
T5 Fully finetuned (no freezing) $0.301\pm 0.007$ $0.448\pm 0.006$ $0.861\pm 0.005$
Freeze first layer $0.350\pm 0.008$ $0.504\pm 0.007$ $0.849\pm 0.006$
Freeze all but final 3 layers $0.367\pm 0.009$ $0.516\pm 0.007$ $0.832\pm 0.006$
Freeze all but final layer $0.598\pm 0.014$ $0.784\pm 0.011$ $0.619\pm 0.011$
Freeze all layers $0.792\pm 0.019$ $0.981\pm 0.013$ $0.420\pm 0.014$
LLaMA-3 Fully finetuned (no freezing) $\mathbf{0.248\pm 0.006}$ $\mathbf{0.345\pm 0.005}$ $\mathbf{0.891\pm 0.004}$
Freeze first layer $0.279\pm 0.007$ $0.426\pm 0.006$ $0.878\pm 0.004$
Freeze all but final 3 layers $0.318\pm 0.008$ $0.474\pm 0.006$ $0.851\pm 0.005$
Freeze all but final layer $0.424\pm 0.010$ $0.576\pm 0.008$ $0.793\pm 0.007$
Freeze all layers $0.716\pm 0.017$ $0.893\pm 0.012$ $0.518\pm 0.013$
@@
In all three models, predictive performance improved progressively as more transformer layers were unfrozen, highlighting the importance of task-specific fine-tuning in scientific regression applications. This trend highlights a key insight: while LLMs capture broadly useful representations during pretraining, these representations alone are insufficient to achieve optimal accuracy in domain-specific prediction tasks without additional adaptation. In the case of RoBERTa and T5, the most notable performance gains occurred when the final three layers of the model were unfrozen. This suggests that the highest-level contextual embeddings learned during pretraining are not directly aligned with the band gap target property but can be effectively adapted through partial retraining.
LLaMA-3, on the other hand, exhibited relatively strong performance even when the majority of its layers were frozen. For example, when only the final three layers were unfrozen during fine-tuning, LLaMA-3 achieved $R^{2}$ values comparable to those of fully fine-tuned RoBERTa and T5 models. This observation points to its efficient architecture and inductive biases, which may make its learned representations more transferable to new scientific tasks. In particular, the decoder-only architecture of LLAMA-3, combined with its use of Rotary Positional Embeddings and SwiGLU activation functions, appears to confer greater flexibility in adapting to structured input formats such as the material strings used in this study. Furthermore, the ability of LLAMA-3 to achieve high $R^{2}$ values with only the final layer trainable supports its utility in resource-constrained scenarios, such as edge inference or few-shot scientific learning contexts.
Collectively, these results reinforce the conclusion that while pretrained LLMs provide a valuable foundation, their effective use in materials science and other physical domains requires carefully calibrated fine-tuning strategies. Strategic layer freezing not only reduces computational burden during training but also offers insight into which parts of the model architecture contribute most directly to scientific prediction tasks. This understanding is critical for efficiently deploying transformer-based models in practical, data-limited, or performance-critical settings.
3.4 Embedding Analysis
@@
We conducted embedding space analysis using t-SNE visualizations to investigate how each model organizes material representations before and after fine-tuning. As shown in Figure 3, we extracted the first-token embeddings from the pretrained and finetuned versions of RoBERTa, T5, and LLaMA-3 using structured string inputs, and colored the points by crystal system. In the “pretrained” condition, which corresponds to the original language model weights before finetuning on the band gap prediction task, the embeddings reflect structural signals learned from general language corpora.
Figure 3: t-sne visualizations of embedding analysis with crystal systems. a pretrained RoBERTa, b finetuned RoBERTa, c pretrained T5, d finetuned T5, e pretrained LLAMA-3, f finetuned LLaMa-3
@@
Figure 4: t-SNE visualizations of embedding analysis for bandgap. a pretrained RoBERTa, b finetuned RoBERTa, c pretrained T5, d finetuned T5, e pretrained LLaMa-3, f finetuned LLaMa-3
@@
In the pretrained state, the embeddings of all three models exhibit meaningful clustering based on the crystal system. This structure-aware behavior is likely driven by explicit features in the input, such as crystal class, space group, and lattice parameters, information that correlates with symmetry. However, while these features can separate materials by structural family, they do not offer sufficient signal to accurately predict electronic properties like band gap. Without finetuning, the models lack the task-specific supervision needed to connect these latent structural cues to quantitative property outcomes. Among the models, LLaMA-3 showed the most distinct clustering by crystal system even before finetuning, particularly in distinguishing high-symmetry groups like cubic and trigonal. RoBERTa and T5 also demonstrate some clustering by structural categories, albeit with greater overlap and more diffuse boundaries. These patterns indicate that pretrained language models are capable of learning latent structural relationships from text, despite not being trained specifically for materials science.
@@
Once finetuned on the band gap prediction task, however, the clustering behavior undergoes a clear transition: embeddings now group primarily by band gap values rather than by crystal system, as visualized in Figure 4. This shift reveals that the fine-tuning process successfully reorients the model’s latent space to emphasize property-relevant dimensions over purely structural ones. The result is an embedding space where materials with similar band gaps are positioned closer together, regardless of their crystallographic classification. This transformation illustrates the effectiveness of supervised training in realigning general-purpose language representations toward targeted scientific objectives.
@@ 4 Conclusion
In this study, we investigated the use of transformer-based language models, RoBERTa, T5, and LLaMA-3, for predicting the band gaps of semiconductor materials directly from textual inputs. We demonstrated that these models can learn meaningful structure–property relationships without relying on handcrafted features or complex numerical descriptors.
Finetuned language models consistently outperformed shallow ML baselines, with LLaMA-3 achieving the highest accuracy using structured inputs (MAE 0.248 eV, $R^{2}$ 0.891). Even with natural language descriptions, the models captured relevant patterns, highlighting their flexibility for scenarios lacking structured data. Additionally, layer freezing analysis showed that strong predictive performance can be achieved with minimal finetuning. These findings underscore the adaptability of language models for scientific regression tasks.
Overall, our results establish transformer-based language models, particularly compact, decoder-only architectures like LLaMA-3, as promising tools for text-driven materials property prediction, offering scalability, interpretability, and efficiency for future materials informatics pipelines.

### v3

@@ 3.2 Model Performance
We evaluated the performance of four transformer-based language models—RoBERTa, T5, Llama-3, and MatSciBERT—on the task of predicting semiconductor band gaps from text-based material descriptions. We also included shallow machine learning baselines using the same input format, including SVR, random forest, and XGBoost. GNNs were not considered, as their graph-based input representations do not align with the feature-based approach used here. Model accuracy was quantified using three metrics: mean absolute error (MAE), root mean square error (RMSE), and the coefficient of determination ($R^{2}$). The results are summarized in Table 2, with parity plots shown in Figure 2.
Table 2: Comparison of model performance across different ML approaches. For LLMs, results are shown as structured string / natural language description. Best performance per metric is shown in bold, and second-best performance is underlined.
@@
MatSciBERT(string) LLM (Encoder) $0.288\pm 0.007$ $0.407\pm 0.006$ $0.871\pm 0.004$
MatSciBERT(description) LLM (Encoder) $0.366\pm 0.009$ $0.503\pm 0.007$ $0.808\pm 0.007$
Figure 2: Parity plots for band gap predictions across models: a SVR, b XGBoost, c Random Forest, d RoBERTa, e T5, f Llama-3, g MatSciBERT
Across all models, inputs in structured string format consistently outperformed descriptive natural language inputs, as shown in Table 2. Both input types are derived from the same underlying features, so the richness of information is equivalent; the difference arises from how the information is presented. Structured strings provide a consistent, systematic representation that reduces variability, making it easier for the model to learn direct mappings between features and target properties. In contrast, language descriptions introduce variability in phrasing, terminology, and length, requiring the model to first infer the relevant features from context. This adds complexity and can obscure patterns, even though the same information is present.
Among the transformer-based models, Llama-3 with structured input achieved the best overall performance, with a mean absolute error of 0.248 eV and $R^{2}$ of 0.891 (Table 2). Following closely, MatSciBERT delivered a strong performance with an MAE of 0.288 eV and an $R^{2}$ of 0.872. While RoBERTa and T5 also outperformed the shallow ML models, their performance was lower when using GPT-generated descriptions compared to structured string inputs.
These results highlight several insights. First, language models can predict band gaps directly from text-based inputs without any feature preprocessing. Given its large number of parameters, it is not surprising that Llama-3 performed the best. More interestingly, when comparing RoBERTa and MatSciBERT, which share similar architectures, MatSciBERT achieved better results despite having fewer parameters. This suggests that the training strategy and the use of a domain-specific corpus are critical. Pretraining on materials science literature enables the model to more effectively capture the relationship between material features and band gap values.
Even though we evaluated multiple types of language models, it is difficult to generalize performance purely based on architecture since the number of parameters varies greatly. Nonetheless, it is important to demonstrate that text-based band gap prediction works across different model types. It is also not surprising that language model approaches outperform shallow machine learning methods. To clarify, both shallow ML models and language models use the same feature set, so the richness of information is identical. The superior performance of language models indicates their enhanced capability to learn complex correlations between the provided features and the band gap.
@@ 3.3 Layer Freezing Analysis
To investigate how pretrained representations in each model contribute to band gap prediction, we conducted layer-freezing experiments on RoBERTa, T5, Llama-3, and MatSciBERT using structured string inputs. In these experiments, we progressively froze layers of the transformer backbone and fine-tuned only a portion of the model, while training a custom regression head on top of the frozen or partially frozen representations. The goal was not to reduce computational cost, but to probe which layers carry the most task-relevant information. Table 3 reports the MAE, RMSE, and $R^{2}$ values for each model under various freezing setups. This analysis provides insight into how much domain-specific fine-tuning is necessary and which parts of the pretrained model are most informative for the downstream task, highlighting the layers that contribute most strongly to predictive accuracy.
Table 3: Comparison of layer freezing strategies across RoBERTa, T5, Llama-3, and MatSciBERT using structured string inputs. The fully finetuned (non-frozen) results are included for reference. Parameter counts and percentages represent the number of trainable parameters relative to the no-freezing case. All values are reported as mean $\pm$ standard deviation.
Model Freezing Strategy Parameters (count, %) MAE (eV) RMSE (eV) $R^{2}$
RoBERTa Fully finetuned (no freezing) 125,236,993 (100%) $0.325\pm 0.006$ $0.447\pm 0.005$ $0.855\pm 0.004$
Freeze first layer 118,149,121 (94.3%) $0.328\pm 0.008$ $0.448\pm 0.006$ $0.848\pm 0.005$
Freeze all but final 3 layers 61,446,145 (49.1%) $0.388\pm 0.009$ $0.510\pm 0.007$ $0.817\pm 0.006$
Freeze all but final layer 8,269,825 (6.6%) $0.509\pm 0.012$ $0.648\pm 0.009$ $0.721\pm 0.009$
Freeze all layers 591,361 (0.5%) $0.663\pm 0.016$ $0.826\pm 0.011$ $0.563\pm 0.013$
T5 Fully finetuned (no freezing) 60,769,793 (100%) $0.301\pm 0.007$ $0.448\pm 0.006$ $0.861\pm 0.005$
Freeze first layer 57,622,785 (94.8%) $0.350\pm 0.008$ $0.504\pm 0.007$ $0.849\pm 0.006$
Freeze all but final 3 layers 21,830,401 (19.8%) $0.367\pm 0.009$ $0.516\pm 0.007$ $0.832\pm 0.006$
Freeze all but final layer 7,671,553 (7.0%) $0.598\pm 0.014$ $0.784\pm 0.011$ $0.619\pm 0.011$
Freeze all layers 592,129 (0.5%) $0.792\pm 0.019$ $0.981\pm 0.013$ $0.420\pm 0.014$
Llama-3 Fully finetuned (no freezing) 1,237,915,649 (100%) $\mathbf{0.248\pm 0.006}$ $\mathbf{0.345\pm 0.005}$ $\mathbf{0.891\pm 0.004}$
Freeze first layer 1,177,094,145 (95.1%) $0.279\pm 0.007$ $0.426\pm 0.006$ $0.878\pm 0.004$
Freeze all but final 3 layers 447,236,097 (36.1%) $0.318\pm 0.008$ $0.474\pm 0.006$ $0.851\pm 0.005$
Freeze all but final layer 325,593,089 (26.3%) $0.424\pm 0.010$ $0.576\pm 0.008$ $0.793\pm 0.007$
Freeze all layers 2,101,249 (0.2%) $0.716\pm 0.017$ $0.893\pm 0.012$ $0.518\pm 0.013$
MatSciBERT Fully finetuned (no freezing) 110,509,825 (100%) $0.288\pm 0.007$ $0.407\pm 0.006$ $0.871\pm 0.004$
Freeze first layer 103,421,953 (93.6%) $0.293\pm 0.007$ $0.404\pm 0.005$ $0.874\pm 0.004$
Freeze all but final 3 layers 22,445,569 (20.3%) $0.340\pm 0.008$ $0.454\pm 0.006$ $0.849\pm 0.005$
Freeze all but final layer 8,269,825 (7.5%) $0.415\pm 0.010$ $0.539\pm 0.007$ $0.798\pm 0.007$
Freeze all layers 1,181,953 (1.1%) $0.827\pm 0.020$ $0.994\pm 0.013$ $0.385\pm 0.015$
@@
In all four models, predictive performance improved progressively as more transformer layers were unfrozen, highlighting the importance of task-specific fine-tuning in scientific regression tasks, as shown in Table 3. While LLMs capture broadly useful representations during pretraining, these representations alone are insufficient to achieve optimal accuracy in domain-specific prediction tasks without additional adaptation. As more layers are involved in fine-tuning, the models can better capture feature–property correlations. Interestingly, MatSciBERT exhibits slightly better RMSE and $R^{2}$ scores compared to its fully finetuned case, which may be attributed to the well-formed representations learned from prior exposure to the materials science domain.
Figure 3: Scaling behavior of finetuning strategies across transformer-based models. MAE is shown as a function of the number of trainable parameters. Colors indicate different model architectures: RoBERTa (yellow), T5 (green), Llama-3 (blue), and MatSciBERT (red). Marker shapes represent different freezing strategies, from fully finetuned (no freezing) to full layer freezing.
The rough scaling relationship between the number of parameters and prediction accuracy is illustrated in Figure 3. Generally, as the number of trainable parameters increases, MAE decreases, indicating improved predictive performance across models and layer-freezing strategies. Within each model, this relationship is relatively linear in the layer-wise freezing experiments, showing that gradually unfreezing layers consistently improves performance. However, when comparing across different models, the relationship is not strictly linear. This reflects the fact that predictive performance depends not only on the number of parameters but also on other factors, including model architecture, pretraining corpus, task-specific fine-tuning strategies, and the alignment between learned representations and the structured input features. These results emphasize that while pretrained LLMs provide a strong foundation, achieving optimal performance in materials science requires careful selection of trainable layers and fine-tuning strategies, especially under computational or resource constraints.
3.4 Feature-wise Self-Attention Score
To identify which material features are most emphasized during band gap prediction, we conducted a feature-wise self-attention analysis on Llama-3 and MatSciBERT, the two best-performing models. In transformer architectures, the attention mechanism quantifies how strongly each token attends to others when forming contextual representations, as shown in Figure 1d.
For each model, input text was tokenized (maximum length of 512 tokens, truncation enabled), and attention weights were extracted from the first and last transformer layers. The attention score here refers to the attention weight between the first token, used as the regression header, and all other tokens in the sequence. For each head, the attention scores from the regression token to the rest of the tokens were averaged, and the resulting values were then averaged across all heads within the layer.
Attention scores between the regression header token and all feature tokens were extracted from each transformer layer. For a given layer $l$ and sample $i$, the attention weights were first averaged across all attention heads:
$s_{i,l}(t)=\frac{1}{H}\sum_{h=1}^{H}s_{i,l,h}(t)$ (1)
where $H$ is the number of attention heads and $s_{i,l,h}(t)$ denotes the attention score from the header token to token $t$ in head $h$.
For each feature $f$, we identified its corresponding token span and recorded the maximum attention score within that span:
$\text{score}_{i,l,f}^{\text{raw}}=\max_{t\in[t_{\text{start}},t_{\text{end}}]}s_{i,l}(t)$ (2)
To ensure comparability across samples, raw scores were normalized to the range [0, 1] using min–max scaling:
$\text{score}_{i,l,f}^{\text{norm}}=\frac{\text{score}_{i,l,f}^{\text{raw}}-\min_{k}\text{score}_{i,l,k}^{\text{raw}}}{\max_{k}\text{score}_{i,l,k}^{\text{raw}}-\min_{k}\text{score}_{i,l,k}^{\text{raw}}}$ (3)
where $k$ indexes all features in sample $i$.
Finally, layer-wise feature attention was obtained by averaging normalized scores across all test samples:
$\text{Avg\_Attention}_{l,f}=\frac{1}{N}\sum_{i=1}^{N}\text{score}_{i,l,f}^{\text{norm}}$ (4)
This procedure provides a consistent measure of how each layer attends to specific material features based on attention scores. It is important to note that these scores do not incorporate the value matrix; therefore, as illustrated in Figure 1d, they do not directly quantify feature importance. Instead, they offer a descriptive view of the model’s focus on the input tokens.
Figure 4: Feature-wise self-attention scores for LLaMA-3 and MatSciBERT. a LLaMA-3, first layer; b LLaMA-3, final layer (layer 16); c MatSciBERT, first layer; d MatSciBERT, final layer (layer 12).
The results reveal clear differences in attention patterns between the first and last layers for both models. In our analysis, the embedding of the first token from the last layer is ultimately used for regression, reflecting how the model allocates attention to optimize prediction performance. For LLaMA-3, attention to geometric features, such as lattice variation, lattice system, and overall geometry, decreases in the final layer, as seen by comparing Figure 4a and b. Although these features are important for structure–property relationships, this pattern suggests that LLaMA-3 struggles to directly extract geometric information from text-based descriptions, instead emphasizing compositional and spin-related features.
MatSciBERT, which benefits from domain-specific pretraining on materials science literature, shows earlier specialization in crystallographic features, as illustrated in Figure 4c. Similar to LLaMA-3, attention to geometric features declines in later layers, highlighting the difficulty of encoding geometric information purely from text (Figure 4d). However, MatSciBERT distributes attention more broadly in the first layer, with structural symmetry attributes such as point group receiving relatively high scores. By the final layer, attention becomes more selective, focusing on key features such as composition, including species and compound information.
3.5 Embedding Map
@@
We conducted embedding space analysis using t-SNE visualizations to investigate how each model organizes material representations before and after fine-tuning. As shown in Figure 5 and 6, we extracted the first-token embeddings from the pretrained and finetuned versions of RoBERTa, T5, Llama-3, and MatSciBERT using structured string inputs, and colored the points by crystal system. In the “pretrained” condition, which corresponds to the original language model weights before finetuning on the band gap prediction task, the embeddings reflect structural signals learned from general language corpora. The “fine-tuned” condition represents embeddings generated after the models were fully fine-tuned for band gap prediction.
Figure 5: t-SNE visualizations of embeddings colored by crystal system. a-d show results from the pretrained models: a RoBERTa, b T5, c Llama-3, d MatSciBERT. e-h show results from the corresponding finetuned models: e RoBERTa, f T5, g Llama-3, h MatSciBERT.
@@
Figure 6: t-SNE visualizations of embeddings for bandgap prediction. a-d show results from the pretrained models: a RoBERTa, b T5, c Llama-3, d MatSciBERT. e-h show results from the corresponding finetuned models: e RoBERTa, f T5, g Llama-3, h MatSciBERT.
@@
In the pretrained state, embeddings from all four models exhibit meaningful clustering according to crystal system, as shown in Figure 5a-d. In contrast, when coloring by bandgap (Figure 6a-d), no clear clustering is observed. This structure-aware behavior likely arises from explicit input features such as crystal class, space group, and lattice parameters, which correlate with symmetry. While these features can separate materials by structural family, they do not provide sufficient signal to predict electronic properties like bandgap. In other words, without finetuning, the models lack task-specific supervision to map structural cues to quantitative property outcomes.
@@
After finetuning on the bandgap prediction task, however, the clustering behavior changes markedly: embeddings now group primarily by bandgap rather than by crystal system, as visualized in Figure 5e-h and Figure 6e-h. In particular, Llama-3 shows a significant loss of crystal system clustering while forming a near-linear embedding map based on bandgap, which likely contributes to its superior predictive performance. The result is an embedding space where materials with similar bandgaps are positioned closer together, independent of their crystallographic classification. This transformation illustrates how supervised training can reorient general-purpose language representations toward property-specific scientific objectives.
T5 produces a qualitatively different t-SNE embedding: unlike the roughly circular maps seen in most other models, T5’s embedding forms a curvy, elongated shape. We note that the models have differing embedding dimensions, and t-SNE can distort distances depending on dimensionality and distribution. Therefore, some apparent differences may reflect visualization artifacts or latent space geometry rather than intrinsic model behavior. Further quantitative analysis of cluster separation and latent structure would be needed to confirm these patterns.
@@ 4 Conclusion
In this study, we investigated the use of transformer-based language models, RoBERTa, T5, Llama-3, and MatSciBERT for predicting the band gaps of semiconductor materials directly from textual inputs. We demonstrated that these models can learn meaningful feature–property relationships without relying on complex feature preprocessing or structure graph of atomic coordinates. Llama-3 achieving the highest accuracy using structured inputs (MAE 0.248 eV, $R^{2}$ 0.891). Even with natural language descriptions, the models captured relevant patterns, highlighting their flexibility for scenarios lacking structured data.
Overall, this work highlights the potential of both general-purpose and domain-specific language models as flexible, scalable, and efficient tools for materials property prediction. By enabling direct prediction of properties from human-readable text, these models remove the need for extensive feature engineering or graph-based structure encoding, allowing rapid, end-to-end property estimation from structured or natural language material descriptions. The results also demonstrate that pretrained language models can provide interpretable insights through attention and embedding analyses, identifying the material features most relevant to the target property. .
@@
Technology Use Disclosure
We used ChatGPT and Claude to help with grammar and typographical corrections during the preparation of this manuscript. The authors have carefully reviewed, verified, and approved all content to ensure accuracy and integrity.